OpenStudy (anonymous):

How to find the length of a hexagon given the apothem 8 square root 3 which is 13.86 in decimal form

OpenStudy (anonymous):

@Rizags

OpenStudy (rizags):



OpenStudy (rizags):

you need side length right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the apothem is 8 square root 3

OpenStudy (rizags):

\[\sin 60 = {\frac{ 8\sqrt{3} }{ x }}\]

OpenStudy (rizags):

so, \[x=\frac{ 8\sqrt{3} }{ \sin 60 }\]

OpenStudy (rizags):

evaluate and tell me ans

OpenStudy (anonymous):

4.453

OpenStudy (rizags):

no, i didnt get that.. try again

OpenStudy (anonymous):

16

OpenStudy (rizags):

yehppppppp plz medal fan again lol

OpenStudy (anonymous):

cool! thanks again and i will